2. As the nurse manager of the new 15-bed observation unit you are responsible for reporting to your nursing director and the finance department any variances identified within your operating budget. Listed below is your monthly operating report. Please review and discuss the specific variances and how you will determine the causes of variances.

Actual

Budget

Variance

Volume/Census

360

420

Supplies

$8,550

$7,800

Nursing Labor

$58,450

$52,575
Answer #1

Census

Variance= 360 - 420 = - 60(unfavourable income).

Supplies

Variance=7,800 - 8,550 = - 750 (unfavourable expenditure).

Nursing Labor

Variance= 52,575 - 58,450 = - 5875 (unfavourable expenditure).

In Census, Supplies and Nursing labor, spenting more than budgeted is evident.So the variances are unfavourable.

Causes of variances

  • Changes in census or not upto expectation
  • Costs of supplies might be higher by the introduction of new technologies
  • More supplies might have used than budgeted.
  • Changes in policies of the unit to use more supplies than before.
  • Nursing salary hike or increment might have lead to unfavourable variance